OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x+y=39 and 3x+4y=136

OpenStudy (anonymous):

x=-y+39 and -4/3y+136/3

OpenStudy (anonymous):

x+y=39 and 3x+4y=136 -y -y -4y -4y x=38 3x= 132 /3 /3 x= 66

OpenStudy (anonymous):

ok thats answer

OpenStudy (anonymous):

kjackson

OpenStudy (anonymous):

that's not on the answer document

OpenStudy (anonymous):

what

OpenStudy (anonymous):

well, looking at my answers, the number could be 10,20,29,or30

OpenStudy (anonymous):

Take the two equations, realize that you can add and subtract the two equations from each other. What I mean by that is, you can add x + y = 39 (from the first equation) to 3x + 4y = 136 (from the second) to get 4x + 5y = 175. Now some cleverness comes into play. We want to get one of the variables to cancel. To do this we can multiply either equation by a constant. Lets multiply the first equation by 4 and subtract it from the second. 4x + 4y - (3x + 4y) = 156 - 136. When you go through and cancel you boil it down to x = 20.

OpenStudy (anonymous):

Oh, okay, that makes sense.

OpenStudy (anonymous):

You can also verify your answer if you need to do this on a test. When you arrive at a conclusion, for instance x = 20. Plug 20 in for x in one of the equations and solve for y. From equation 1 we get 20 + y = 39, which we can see y = 19. Now plugging that into equation 2 we get 3(20) + 4(19) = 136 so our answer is correct.

OpenStudy (anonymous):

Oh, thanks! I understand.

OpenStudy (anonymous):

mary100 your answer is wrong there is only one answer not to. The correct answer is x=20 it is that answer because you can solve equations by graphing x+y=39 3x+4y=136 Put your answer -x -x -3x -3x in your calc and graph y=-x+39 4y=-3x+136 it than you will see 20 divide by 4 than check your y=-3/4x+34 answer

OpenStudy (anonymous):

so I can do this with or without a calculator?

OpenStudy (anonymous):

You can do it without a calculator but the calculator just checks it

OpenStudy (anonymous):

Coming from someone with a B.S. in Math, you NEVER NEED a calculator, but they are often helpful.

OpenStudy (anonymous):

Well, I came from poor schools and back then we never got to use calculators.

OpenStudy (anonymous):

If you do solving equations by graphing you need a calculator but if you do elimination or substitution you totally do not need a calculator

OpenStudy (anonymous):

thanks, you are great!!